OpenStudy (anonymous):

need help with finding the domain of a function.. g(x)=x+7 over x^2-16

OpenStudy (anonymous):

x cannot = 4

OpenStudy (anonymous):

or -4

OpenStudy (anonymous):

be careful here. you need to solve \[x^2-16=0\] and there are two solutions, namely 4 and -4

OpenStudy (anonymous):

x can't be 4 or -4

OpenStudy (anonymous):

I forgot the +/- symbol

OpenStudy (anonymous):

the answer has to be expressd in interval notation

OpenStudy (anonymous):

(infinity, -4) (-4,4)(4, infinity)

OpenStudy (anonymous):

thank you for your help!

OpenStudy (anonymous):

Or like this: \[x \in (-\infty,-4) \cup (-4,4) \cup (4,\infty)\]

OpenStudy (anonymous):

ooh that looks nice!

OpenStudy (anonymous):

is that \[x\in \]

OpenStudy (anonymous):

got it on the first try